A passenger train's speed is 60mi/h,and a freight train's is 40mi/h. The passenger train travels the same distance in 1.5h less time than the freight train. How long does each train take to make the trip?
Answer by ankor@dixie-net.com(22740)

A passenger train's speed is 60mi/h, and a freight train's is 40mi/h.
The passenger train travels the same distance in 1.5h less time than the freight train.
How long does each train take to make the trip?
:
let t = pass train travel time
then
(t+1.5) = freight's time
:
Write a distance equation: dist = speed * time
:
Pass dist = freight dist
60t = 40(t+1.5)
60t = 40t + 60
69t - 40t = 60
20t = 60
t = 60/20
t = 3 hrs for pass train to make the trip
then
3 + 1.5 = 4.5 hrs for freight to do it
:
:
Check the solution, by finding the dist each traveled, should be equal
3(60) = 180 mi
4.5(40)= 180 mi